OpenStudy (anonymous):

Given that A=84, B=66 and a=13 solve triangle ABC. Round to the nearest hundredth.

OpenStudy (mathmale):

Since an angle and the side opposite are given (84 degrees, 13), and we know one more angle (B: 66 degrees), we can use the Law of Sines to find side b. Try the following:\[\frac{ b }{ \sin B }=\frac{ a }{ \sin A }\]

OpenStudy (anonymous):

?